One person produces heat at a rate of 5454 W. What is the temperature rise produced in 3030 min by two people in an airtight room of 4747 m3 volume, initially at 0degrees Upper C ° C and at 105 Pa? It is supposed that all the heat is absorbed by the air. Upper K right parenthesis = 0.72 kJ / (kg · K) and M = 29 g / mol are given
1
Expert's answer
2020-09-14T10:20:16-0400
"V=const\to \Delta Q=mc_V\Delta T\\Pt=\frac{pVM}{RT}c_V\Delta T\\(54)(30\cdot 60)=\frac{(10^5)(47)(0.029)}{8.31\cdot 273}720\Delta T\\\Delta T=2.25\degree C"